# Integrated Push-and-Pull Update Model for Goal-Oriented Effective Communication

### v1

@@ III-A2 Action lateness
The lateness of an action performed in the $n$-th time slot in relevance to a query raised at the $n^{\prime}$-th slot, $n^{\prime}\leq n$, is given by
@@
which is valid for $\Theta_{n}<\Theta_{\rm max}$. Herein, $\Theta_{\rm max}$ shows the width of action window within which the AA can act on each query based on update arrivals from the SA. Outside the dedicated action window for the SA, the AA might undertake other tasks or communicate with other agents. Employing the push-based, pull-based, and push-and-pull update communication models, we have $\Theta_{\rm max}=\infty$, $\Theta_{\rm max}=1$, and $\Theta_{\rm max}>1$, respectively. Fig. 3 illustrates the action and idle windows for different models. It is worth mentioning that a wider action window allows for higher flexibility in cases of heavy action loads at the cost of longer actuation availability.
[…]
[scale=0.86]Media/Query_Table.eps
@@ III-C Effectiveness Indicator
An update in the $n$-th time slot is considered effective at the system level if its $\operatorname{GoE}_{n}$ is higher than a target effectiveness grade, which is called $\operatorname{GoE}_{\rm tgt}$ and is necessary to satisfy the goal. Let us define $E_{n}$ as an effectiveness indicator in the $n$-th time slot. Thus, we can write
@@
The second condition in (4) appears from (3). According to (4), an update could be effective only if it arrives within the action window of the AA. Hence, a consequent E-ACK shared with the SA can imply the raise of a query or the availability of the AA to take action. Given the values of $\Delta_{n}$ and $\Theta_{n}$, and by inserting (1) into (4), we reach a target usefulness level $v_{\rm tgt}$ as the importance threshold that the update should exceed to be considered effective. In this case, if $\Theta_{n}<\Theta_{\rm max}$, we have
@@
$\operatorname{min}\!\left\{\hat{\nu}_{j}~{}|~{}\hat{\nu}_{j}\in\hat{\mathcal{V}},f_{g}(g_{\Delta}(\hat{\nu}_{j},\Delta_{n}),g_{\Theta}(\Theta_{n}))\geq\operatorname{GoE}_{\rm tgt}\!\right\}\!;$ (5)
otherwise, $v_{\rm tgt}=\operatorname{max}\{\hat{\nu}_{j}~{}|~{}\hat{\nu}_{j}\in\hat{\mathcal{V}}\}$. In (III-C), $v_{\rm tgt}$ can be computed by exhaustive search.
@@ IV Model-Based Agent Decisions
In this section, we first formulate a decision problem for effect-aware policies at either or both agent(s), cast it as a constrained Markov decision process (CMDP), and then solve it based on the problem’s dual form.
@@ IV-A Problem Formulation
The objective is to maximize the expected discounted sum of the updates’ effectiveness in fulfilling the subscribed goal, where each agent individually derives its decision policy subject to the relevant ensued cost by looking into the problem from its own perspective. Let us define $\pi_{\alpha}^{*}$ and $\pi_{\beta}^{*}$ as the classes of optimal policies for transmission and query controls, respectively. Therefore, we can formulate the decision problem solved at each agent as follows
@@ IV-B1 Modeling at the SA
States – The state of the system $S_{\alpha,n}$ in the $n$-th slot from the SA’s perspective is depicted by a tuple $(v_{n},\hat{E}_{n})$ in which $v_{n}$ is the update’s usefulness, and $\hat{E}_{n}\in\{0,1\}$ shows the E-ACK arrival status at the SA after passing the PEC, as defined in Section II. Herein, we have $\hat{E}_{n}=0$ in case $E_{n}=0$ or the acknowledgment signal is erased; otherwise, $\hat{E}_{n}=1$. In this regard, $S_{\alpha,n}$ belongs to a finite and countable state space $\mathcal{S}_{\alpha}$ with $\lvert\mathcal{S}_{\alpha}\rvert=2\cdot\lvert\mathcal{V}\rvert$ elements.
Actions – We consider $\alpha_{n}$ the decision for update communication in the $n$-th slot, which is a member of an action space $\mathcal{A}_{\alpha}=\{0,1\}$. In this space, 00 stands for discarding the update, and $1$ indicates transmitting the update.
@@
$p_{\alpha}(S_{\alpha,n},\alpha_{n},S_{\alpha,n+1})$ $=\mathrm{Pr}\big{(}(v_{n+1},\hat{E}_{n+1})~{}|~{}(v_{n},\hat{E}_{n}),\alpha_{n}\big{)}$
$=p_{\nu}(v_{n+1})\mathrm{Pr}\big{(}\hat{E}_{n+1}~{}|~{}v_{n},\alpha_{n}\big{)}$ (7)
@@
  • $\mathrm{Pr}\big{(}\hat{E}_{n+1}=0~{}|~{}v_{n},\alpha_{n}\big{)}=\mathrm{Pr}(\hat{v}_{\rm tgt}>\alpha_{n}v_{n})=1-P_{\hat{v}_{\rm tgt}}(\alpha_{n}v_{n})$,
  • $\mathrm{Pr}\big{(}\hat{E}_{n+1}=1~{}|~{}v_{n},\alpha_{n}\big{)}=P_{\hat{v}_{\rm tgt}}(\alpha_{n}v_{n})$,
where $\hat{v}_{\rm tgt}$ is a mapped target usefulness that the SA considers, $P_{\hat{v}_{\rm tgt}}(\hat{v}_{\rm tgt})=\sum_{\hat{v}_{\rm tgt}^{\prime}\leq\hat{v}_{\rm tgt}}p_{\hat{v}_{\rm tgt}}(\hat{v}_{\rm tgt}^{\prime})$ denotes its cumulative distribution function (CDF), and $p_{\hat{v}_{\rm tgt}}(\cdot)$ shows the pmf derived in Section V-B.
@@ IV-B2 Modeling at the AA
For modeling the problem at the AA, we have the components as follows: States – We represent the state $S_{\beta,n}$ in the $n$-th time slot using a tuple $(\hat{v}_{n},\Delta_{n},\Theta_{n})$, where $\hat{v}_{n}$ is the usefulness of the received update from the perspective of the endpoint, $\Delta_{n}$ is the AoI, and $\Theta_{n}$ denotes the action lateness, as modeled in Section III. Without loss of generality, we assume the values of $\Delta_{n}$ and $\Theta_{n}$ are truncated by the maximum values notated as $\Delta_{\rm max}$ and $\Theta_{\rm max}$, respectively, such that the conditions
@@
Actions – As already mentioned, $\beta_{n}$ shows the decision of raising a query in the $n$-th time slot and gets values from an action space $\mathcal{A}_{\beta}=\{0,1\}$. Here, 00 and $1$ depict refusing and confirming to pull an update, respectively.
@@
$~{}~{}~{}~{}~{}~{}~{}\mathrm{Pr}\big{(}(\hat{v}_{n+1},\Delta_{n+1},\Theta_{n+1})~{}|~{}(\hat{v}_{n},\Delta_{n},\Theta_{n}),\beta_{n}\big{)}.$ (10)
@@
  • $\mathrm{Pr}((\hat{\nu}_{j},\operatorname{min}\{\Delta_{n}\!+\!1,\Delta_{\rm max}\},\\ \operatorname{min}\{\Theta_{n}\!+\!1,\Theta_{\rm max}\})~{}|~{}(\hat{\nu}_{j},\Delta_{n},,\Theta_{n}),\beta_{n})=1-\beta_{n}$,
  • $\mathrm{Pr}((\hat{\nu}_{j},\operatorname{min}\{\Delta_{n}\!+\!1,\Delta_{\rm max}\},1)~{}|~{}(\hat{\nu}_{j},\Delta_{n},\Theta_{n}),\beta_{n})=\beta_{n}p_{\epsilon}$,
  • $\mathrm{Pr}((\hat{\nu}_{j^{\prime}},1,1)~{}|~{}(\hat{\nu}_{j},\Delta_{n},\Theta_{n}),\beta_{n})=\beta_{n}(1\!-\!p_{\epsilon})q_{j^{\prime}}$,
@@ IV-B3 Independence of the initial state
Given the transition probabilities defined in (IV-B1), every state $S_{\alpha,m}\in\mathcal{S}_{\alpha}$, $m\leq N$, is accessible or reachable from the state $S_{\alpha,n}$ in finite steps with a non-zero probability, following the policy $\pi_{\alpha}$. Therefore, the accessibility condition holds for the CMDP model at the SA [21, Definition 4.2.1]. ∎
@@
We divide the state space $\mathcal{S}_{\beta}$ into two disjoint spaces of $\mathcal{T}_{a}$ and $\mathcal{T}_{b}=\mathcal{S}_{\beta}-\mathcal{T}_{a}$, where $\mathcal{T}_{a}$ consists of all the states whose $\Delta_{n}=1$, i.e., $\mathcal{T}_{a}=\{S_{\beta,n}~{}|~{}S_{\beta,n}=({\hat{\nu}}_{j},1,\Theta_{n}),\forall\hat{\nu}_{j}\in\hat{\mathcal{V}},\Theta_{n}=1,2,...,\Theta_{\rm max}\}$. Thus, $\mathcal{T}_{b}$ includes the rest of the states with $\Delta_{n}\geq 2$. With regard to the transition probabilities derived in (IV-B2), all states of $\mathcal{T}_{b}$ are transient under any policy, while every state of an arbitrary pair of two states in $\mathcal{T}_{a}$ is accessible from the other state. Accordingly, the weak accessibility condition in the modeled CMDP at the AA is satisfied according to [21, Definition 4.2.2]. ∎
@@
$\mathcal{P}_{2}:~{}$ $\underset{\pi_{\gamma}}{\operatorname{max}}~{}\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}E_{n}\bigg{]}\coloneqq\bar{E}_{\gamma}$
@@ IV-C Dual Problem
$\mathcal{L}(\mu;\pi_{\gamma})$ $=\underset{\pi_{\gamma}}{\operatorname{max}}~{}\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}\Big{(}E_{n}-\mu c_{\gamma}(\gamma_{n})\Big{)}\bigg{]}$
@@
for $\mu$ derived in the dual problem $\mathcal{P}_{3}$. As the dimension of the state space $\mathcal{S}_{\gamma}$ is finite for both defined models, the growth condition is met [22]. Also, the immediate reward is bounded below, having a non-negative value according to Section IV-B. In light of the above satisfied conditions, from [22, Corollary 12.2], we can claim that $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ converge to the same expected values, thus we have
$\bar{E}_{\gamma}=\underset{\mu\geq 0}{\inf}~{}h_{\gamma}(\mu)=\underset{\pi_{\gamma}}{\operatorname{max}}~{}u(\pi_{\gamma})$ (16)
under any class of policy $\pi_{\gamma}$. Owing to the satisfied conditions, there exist non-negative optimal values for the Lagrange multiplier $\mu^{*}$ such that we can define $u(\pi_{\gamma})=\mathcal{L}(\mu^{*};\pi_{\gamma})$ in (16) [22, Theorem 12.8].
@@ IV-D Iterative Algorithm
Input: Known parameters $N\gg 1$, $C_{\gamma,\rm max}$, $\eta$, $\varepsilon_{\mu}$, state space $\mathcal{S}_{\gamma}$, and action space $\mathcal{A}_{\gamma}$. The form of the cost function $c_{\gamma}(\cdot)$. Initial values $l\leftarrow 1$, $\mu^{(0)}\leftarrow 0$, $\mu^{-}\leftarrow 0$, $\mu^{+}\gg 1$, $\pi_{\gamma,\mu^{-}}\leftarrow 0$, and $\pi_{\gamma,\mu^{+}}\leftarrow 0$.
1 Initialize $\pi_{\gamma,\mu}^{*}(s)$, $\forall s\in\mathcal{S}_{\gamma}$, via running Utility($\mu^{(0)}$).
@@
4       update $\mu^{(l)}\leftarrow\frac{\mu^{-}+\mu^{+}}{2}$.
5       Improve $\pi_{\gamma,\mu}^{*}\leftarrow$ Utility($\mu^{(l)}$).
@@
7            $\mu^{-}\leftarrow\mu^{(l)}$, and $\pi_{\gamma,\mu^{-}}\leftarrow$ Utility($\mu^{-}$).
8      else $\mu^{+}\leftarrow\mu^{(l)}$, and $\pi_{\gamma,\mu^{+}}\leftarrow$ Utility($\mu^{+}$).
9       Reset $l\leftarrow l+1$.
10if $\mathbb{E}\Big{[}\sum_{n=1}^{N}c_{\gamma}(\gamma_{n})\Big{]}<NC_{\gamma,\rm max}$ then $\pi_{\gamma,\mu}^{*}(s)\leftarrow\eta\pi_{\gamma,\mu^{-}}(s)+(1-\eta)\pi_{\gamma,\mu^{+}}(s)$, $\forall s\in\mathcal{S}_{\gamma}$.
@@
Function Utility($\mu$):
       Input: Known parameters $N\gg 1$, $\varepsilon_{\pi}$, state space $\mathcal{S}_{\gamma}$, and action space $\mathcal{A}_{\gamma}$. Initial values $k\leftarrow 1$, $\pi_{\gamma,\mu}(s)\leftarrow 0$, and $V_{k}^{\pi_{\gamma,\mu}}(s)\leftarrow 0$, $\forall s\in\mathcal{S}_{\gamma}$.

### v2

@@ III-A2 Action lateness
The lateness of an action performed at the $n$-th time slot in relevance to a query initiated at the $n^{\prime}$-th slot, $n^{\prime}\leq n$, is calculated as follows
@@
which is valid for $\Theta_{n}<\Theta_{\rm max}$. Herein, $\Theta_{\rm max}$ shows the width of action window within which the AA can act on each query based on update arrivals from the SA. Outside the dedicated action window for the SA, the AA might undertake other tasks or communicate with other agents. Employing the push-based, pull-based, and push-and-pull update communication models, we have $\Theta_{\rm max}=\infty$, $\Theta_{\rm max}=1$, and $\Theta_{\rm max}>1$, respectively. Fig. 2 shows the action and idle windows for different models. A wider action window enables higher flexibility during heavy action loads at the cost of longer actuation availability.
@@ III-C Effectiveness Indicator
An update at the $n$-th time slot is considered effective at the system level if its $\operatorname{GoE}_{n}$ is higher than a target effectiveness grade, which is called $\operatorname{GoE}_{\rm tgt}$ and is necessary to satisfy the goal. Let us define $E_{n}$ as an effectiveness indicator at the $n$-th time slot. Thus, we can write
@@
The second condition in (4) appears from (3). According to (4), an update could be effective only if it arrives within the action window of the AA. Hence, a consequent E-ACK shared with the SA can imply the initiation of a query or the availability of the AA to take action. Given the values of $\Delta_{n}$ and $\Theta_{n}$, and by inserting (1) into (4), we reach a target usefulness level $v_{\rm tgt}$ as the importance threshold that the update should exceed to be considered effective. In this case, if $\Theta_{n}<\Theta_{\rm max}$, we have
@@
$\operatorname{min}\left\{\hat{\nu}_{j}\,\lvert\,\hat{\nu}_{j}\in\hat{\mathcal{V}},f_{g}(g_{\Delta}(\hat{\nu}_{j},\Delta_{n}),g_{\Theta}(\Theta_{n}))\geq\operatorname{GoE}_{\rm tgt}\right\}\!;$ (5)
otherwise, $v_{\rm tgt}=\operatorname{max}\{\hat{\nu}_{j}\,\lvert\,\hat{\nu}_{j}\in\hat{\mathcal{V}}\}$. In (III-C), $v_{\rm tgt}$ can be computed by exhaustive search.
@@ IV Model-Based Agent Decisions
Here, we first formulate a decision problem for effect-aware policies, cast it as a constrained Markov decision process (CMDP), and then solve it based on the problem’s dual form.
@@ IV-A Problem Formulation
The objective is to maximize the expected discounted sum of the updates’ effectiveness in fulfilling the subscribed goal, where each agent individually derives its decision policy subject to the relevant ensued cost by looking into the problem from its own perspective. Let us define $\pi_{\alpha}^{*}$ and $\pi_{\beta}^{*}$ as the classes of optimal policies for transmission and query controls, respectively. Therefore, we can formulate the decision problem solved at each agent as follows [^4] In an ideal scenario, where both agents have full knowledge of the goals and the source’s evolution, the problem could be approached in a centralized manner, allowing for the joint derivation of policies for both agents. However, this falls outside the scope of this work.
@@ IV-B1 Modeling at the SA
States – The state of the system $S_{\alpha,n}$ at the $n$-th slot from the SA’s perspective is depicted by a tuple $(v_{n},\hat{E}_{n})$ in which $v_{n}$ is the update’s usefulness, and $\hat{E}_{n}\in\{0,1\}$ shows the E-ACK arrival status at the SA after passing the PEC, as defined in Section II. Herein, we have $\hat{E}_{n}=0$ in case $E_{n}=0$ or the acknowledgment signal is erased; otherwise, $\hat{E}_{n}=1$. In this regard, $S_{\alpha,n}$ belongs to a finite and countable state space $\mathcal{S}_{\alpha}$ with $\lvert\mathcal{S}_{\alpha}\rvert=2\cdot\lvert\mathcal{V}\rvert$ elements.
Actions – We denote $\alpha_{n}$ the decision for update communication at the $n$-th slot, which is a member of an action space $\mathcal{A}_{\alpha}=\{0,1\}$. In this space, 00 stands for discarding the update, and $1$ indicates transmitting the update.
@@
$p_{\alpha}(S_{\alpha,n},\alpha_{n},S_{\alpha,n+1})$ $=\mathrm{Pr}\big{(}(v_{n+1},\hat{E}_{n+1})\,\lvert\,(v_{n},\hat{E}_{n}),\alpha_{n}\big{)}$
$=p_{\nu}(v_{n+1})\mathrm{Pr}\big{(}\hat{E}_{n+1}\,\lvert\,v_{n},\alpha_{n}\big{)}$ (7)
@@
  • $\mathrm{Pr}\big{(}\hat{E}_{n+1}=0\,\lvert\,v_{n},\alpha_{n}\big{)}=\mathrm{Pr}(\hat{v}_{\rm tgt}>\alpha_{n}v_{n})=1-P_{\hat{v}_{\rm tgt}}(\alpha_{n}v_{n})$,
  • $\mathrm{Pr}\big{(}\hat{E}_{n+1}=1\,\lvert\,v_{n},\alpha_{n}\big{)}=P_{\hat{v}_{\rm tgt}}(\alpha_{n}v_{n})$,
where $\hat{v}_{\rm tgt}$ is a mapped target usefulness that the SA considers, $P_{\hat{v}_{\rm tgt}}(\hat{v}_{\rm tgt})=\sum_{\hat{v}_{\rm tgt}^{\prime}\leq\hat{v}_{\rm tgt}}p_{\hat{v}_{\rm tgt}}(\hat{v}_{\rm tgt}^{\prime})$ indicates its cumulative distribution function (CDF), and $p_{\hat{v}_{\rm tgt}}(\cdot)$ shows the pmf derived in Section V-B.
@@ IV-B2 Modeling at the AA
For modeling the problem at the AA, we have the components as follows:
States – We represent the state $S_{\beta,n}$ at the $n$-th time slot using a tuple $(\hat{v}_{n},\Delta_{n},\Theta_{n})$, where $\hat{v}_{n}$ is the usefulness of the received update from the perspective of the endpoint, $\Delta_{n}$ is the AoI, and $\Theta_{n}$ denotes the action lateness, as modeled in Section III. Without loss of generality, we assume the values of $\Delta_{n}$ and $\Theta_{n}$ are truncated by the maximum values notated as $\Delta_{\rm max}$ and $\Theta_{\rm max}$, respectively, such that the conditions
@@
Actions – As already mentioned, $\beta_{n}$ shows the decision of raising a query at the $n$-th time slot and gets values from an action space $\mathcal{A}_{\beta}=\{0,1\}$. Here, 00 and $1$ depict refusing and confirming to pull an update, respectively.
@@
$~{}~{}~{}~{}~{}~{}~{}\mathrm{Pr}\big{(}(\hat{v}_{n+1},\Delta_{n+1},\Theta_{n+1})\,\lvert\,(\hat{v}_{n},\Delta_{n},\Theta_{n}),\beta_{n}\big{)}.$ (10)
@@
  • $\mathrm{Pr}((\hat{\nu}_{j},\operatorname{min}\{\Delta_{n}\!+\!1,\Delta_{\rm max}\},\\ \operatorname{min}\{\Theta_{n}\!+\!1,\Theta_{\rm max}\})\,\lvert\,(\hat{\nu}_{j},\Delta_{n},,\Theta_{n}),\beta_{n})=1-\beta_{n}$,
  • $\mathrm{Pr}((\hat{\nu}_{j},\operatorname{min}\{\Delta_{n}\!+\!1,\Delta_{\rm max}\},1)\,\lvert\,(\hat{\nu}_{j},\Delta_{n},\Theta_{n}),\beta_{n})=\beta_{n}p_{\epsilon}$,
  • $\mathrm{Pr}((\hat{\nu}_{j^{\prime}},1,1)\,\lvert\,(\hat{\nu}_{j},\Delta_{n},\Theta_{n}),\beta_{n})=\beta_{n}(1\!-\!p_{\epsilon})q_{j^{\prime}}$,
@@ IV-B3 Independence of the initial state
Given the transition probabilities defined in (IV-B1), every state $S_{\alpha,m}\in\mathcal{S}_{\alpha}$, $m\leq N$, is accessible or reachable from the state $S_{\alpha,n}$ in finite steps with a non-zero probability, following the policy $\pi_{\alpha}$. Therefore, the accessibility condition holds for the CMDP modeled at the SA [21, Definition 4.2.1]. ∎
@@
We divide the state space $\mathcal{S}_{\beta}$ into two disjoint spaces of $\mathcal{T}_{a}$ and $\mathcal{T}_{b}=\mathcal{S}_{\beta}-\mathcal{T}_{a}$, where $\mathcal{T}_{a}$ consists of all the states whose $\Delta_{n}=1$, i.e., $\mathcal{T}_{a}=\{S_{\beta,n}\,\lvert\,S_{\beta,n}=({\hat{\nu}}_{j},1,\Theta_{n}),\forall\hat{\nu}_{j}\in\hat{\mathcal{V}},\Theta_{n}=1,2,\dots,\Theta_{\rm max}\}$. Thus, $\mathcal{T}_{b}$ includes the rest of the states with $\Delta_{n}\geq 2$. With regard to the transition probabilities derived in (IV-B2), all states of $\mathcal{T}_{b}$ are transient under any policy, while every state of an arbitrary pair of two states in $\mathcal{T}_{a}$ is accessible from the other state. Accordingly, the weak accessibility condition in the modeled CMDP at the AA is satisfied according to [21, Definition 4.2.2]. ∎
@@
$\mathcal{P}_{2}:~{}$ $\underset{\pi_{\gamma}}{\operatorname{max}}~{}\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}E_{n}\bigg{]}$
@@ IV-C Dual Problem
$\mathcal{L}(\mu;\pi_{\gamma})$ $=~{}\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}\Big{(}E_{n}-\mu c_{\gamma}(\gamma_{n})\Big{)}\bigg{]}$
@@
for $\mu$ derived in the dual problem $\mathcal{P}_{3}$. As the dimension of the state space $\mathcal{S}_{\gamma}$ is finite for both defined models, the growth condition is met [22]. Furthermore, the immediate reward and the induced cost are bounded according to Section IV. Given these satisfied conditions, from [22, Corollary 12.2], we can assert that $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ converge to the same expected values such that we can write
$\underset{\mu\geq 0}{\inf}~{}\underset{\pi_{\gamma}}{\operatorname{max}}~{}\mathcal{L}(\mu;\pi_{\gamma})=\underset{\mu\geq 0}{\inf}~{}\mathcal{L}(\mu;\pi_{\gamma}^{*})=\underset{\pi_{\gamma}}{\operatorname{max}}~{}\mathcal{L}(\mu^{*};\pi_{\gamma})$ (16)
for some policy class $\pi_{\gamma}^{*}$. Given the conditions satisfied, there exist non-negative optimal values for the Lagrange multiplier $\mu^{*}$ under Slater’s condition such that (16) holds [22, Theorem 12.8], establishing the strong duality between $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$.
@@ IV-D Iterative Algorithm
Input: Given parameters $N\gg 1$, $C_{\gamma,\rm max}$, $\eta$, $\varepsilon_{\mu}$, CMDP’s state space, i.e., $\mathcal{S}_{\gamma}$, and action space, i.e., $\mathcal{A}_{\gamma}$. The form of the cost function $c_{\gamma}(\cdot)$. Initial values $l\leftarrow 0$, $\mu^{(0)}\leftarrow 0$, $\mu_{-}^{(0)}\leftarrow 0$, $\mu_{+}^{(0)}>0$, $\pi_{\gamma,\mu_{-}^{(0)}}\leftarrow 0$, and $\pi_{\gamma,\mu_{+}^{(0)}}\leftarrow 0$.
1 Derive $\pi_{\gamma,\mu}^{*}(s)$, $\forall s\in\mathcal{S}_{\gamma}$, via running policy($\mu^{(0)}$).
@@
4       set $l\leftarrow l+1$, $\mu_{-}^{(l)}\leftarrow\mu_{-}^{(l-1)}$, and $\mu_{+}^{(l)}\leftarrow\mu_{+}^{(l-1)}$.
5       Update $\mu^{(l)}\leftarrow\frac{\mu_{-}^{(l)}+\mu_{+}^{(l)}}{2}$.
6       Improve $\pi_{\gamma,\mu}^{*}\leftarrow$ policy($\mu^{(l)}$).
@@
8            $\mu_{-}^{(l)}\leftarrow\mu^{(l)}$, and $\pi_{\gamma,\mu_{-}^{(l)}}\leftarrow$ policy($\mu_{-}^{(l)}$).
9      else $\mu_{+}^{(l)}\leftarrow\mu^{(l)}$, and $\pi_{\gamma,\mu_{+}^{(l)}}\leftarrow$ policy($\mu_{+}^{(l)}$).
10
11if $\mathbb{E}\Big{[}\sum_{n=1}^{N}c_{\gamma}(\gamma_{n})\Big{]}<NC_{\gamma,\rm max}$ then $\pi_{\gamma,\mu}^{*}(s)\leftarrow\eta\pi_{\gamma,\mu_{-}^{(l)}}(s)+(1-\eta)\pi_{\gamma,\mu_{+}^{(l)}}(s)$, $\forall s\in\mathcal{S}_{\gamma}$.
@@
Function policy($\mu$):
       Input: Known parameters from the outer loop. Initial values $k\leftarrow 1$, $\pi_{\gamma,\mu}(s)\leftarrow 0$, and $V_{k}^{\pi_{\gamma,\mu}}(s)\leftarrow 0$, $\forall s\in\mathcal{S}_{\gamma}$.